I am so lost with this graphing stuff, can you please help me with this one?
2x + 3y = 9, I came up with 2(3) + 3(1) = 9 (6,3) is this right, and are there other coordinates? I can't seem to come up with any others.

Answer by wowkids38(4)

Simplifying
2x + 3y = 9
Solving
2x + 3y = 9
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Add '-3y' to each side of the equation.
2x + 3y + -3y = 9 + -3y
Combine like terms: 3y + -3y = 0
2x + 0 = 9 + -3y
2x = 9 + -3y
Divide each side by '2'.
x = 4.5 + -1.5y
Simplifying
x = 4.5 + -1.5y